# Enhancing LLM-based Code Translation in Repository Context via Triple Knowledge-Augmented

### v2

@@ 2.4. Knowledge-Augmented Code Translation
LLM-based code repair. The code translated by LLMs may contain simple syntax errors (e.g., performing mutable operations on immutable variables). Previous work (Deligiannis et al., 2024) has shown that LLMs can repair certain errors in code using compiler error messages. After translation, K3Trans provides the LLM with specific instructions to leverage error message and iteratively refine translation results that fail testing due to compilation or functional errors. While this process can be conducted for multiple iterations, our implementation limits it to one iteration for cost-effectiveness. This approach further improves translation accuracy and assesses the percentage of erroneous translations that can be successfully repaired. The prompt used for code repair is shown as follows:
You were asked to translate the given {corpus_lang} function to {query_lang} according to the {query_lang} function signature, {query_lang} function dependencies (including function and variable dependencies), and data type declarations and {query_lang} function dependency libraries I provide (delimited with XML tags). Some errors occurred when executing your code. Fix the error in your previous response. Only response the function results. ## Basic Details about the function (the same as Translation’s Required Information in translation process) ……
@@ 3. Experiment Setup
  • RQ5 (Bad Case Study): Why does K3Trans fail in translating some codes?
@@ 3.1. Benchmark
We choose RustRepoTrans (Ou et al., 2024) for the following reasons. RustRepoTrans is the first repository-level context code translation benchmark, constructed from GitHub projects that reflect the complexities of real-world software development. As shown in previous work (Ou et al., 2024), LLMs exhibit much poorer performance on RustRepoTrans than on other benchmarks due to the repository-level dependencies and context. This shortcoming of LLM in dealing with real translation scenarios is exactly what K3Trans aims to solve.
Studied Programming Languages. We focus on repository-level code translation for translating code from C, Java and Python to Rust, which are the same supported language pairs in RustRepoTrans. A more important reason is that Rust, as the target language, is closer to the actual requirements of real-world code translation scenarios (Security, 2024). Furthermore, without language-specific adaptation, if K3Trans performs well with a low-resource language such as Rust, it can be reasonably inferred that K3Trans would also be applicable to other high-resource languages.
@@ 3.2. Baselines
We evaluates the performance of K3Trans with two baselines: LLM-based with prompt engineering (PE) and Basic RAG.
@@ 3.3. Metrics
  • Matchast: The Matchast is part of CodeBLEU, the reason we extracted it separately is to evaluate the ability of LLMs to recognize syntax differences. If the model can accurately identify syntax differences and avoid generating corresponding check statements (e.g. if statements) in the target language, the paths of generated code are more similar to the ground truth, thus resulting in a higher Matchast.
Execution-based Metrics. The execution-based metrics aim to assess the quality of generated code through the execution status of the code. We employed Compilation@k, Pass@k, and DSR@k for evaluation:
@@
[…]
@@
[…]
  • Repairable Ratio: Repairable Ratio (RR), as expressed in Equation 6, is an evaluation metric for code translation tasks that measures the proportion of incorrect translations that can be automatically repaired using a LLM. This metric reflects the model’s ability to generate code that, while initially incorrect, is close enough to a correct implementation that it can be fixed with minimal intervention.
    (6) RR=DSR@k - Pass@k1 - $\text{RR}=\frac{\text{DSR@k - Pass@k}}{\text{1 - Pass@k}}$
@@ 3.4. Implementations
We construct the 100% state of the knowledge base, where all functions except the target function have been successfully translated and can serve as sources for dependency usage examples and translation function pair examples, along with all currently available GitHub open-source repositories to build target-language code samples. This setup, referred to as K3Trans in this paper, is used to evaluate the theoretical best performance that an LLM enhanced with self-evolving RAG can achieve in code translation tasks. Additionally, for online code translation, we use the Claude-3.5-Sonnet model (Anthropic, 2024) as it outperformed the other LLMs in repository-level context code translation (Ou et al., 2024). For the knowledge retrieval process, we utilize Elasticsearch (Elasticsearch, 2025) as our search engine, which based on the Lucene library using BM25 as the default score function.
@@ 4.1. RQ1: Performance of K3Trans
Table 1. Comparison of K3Trans and Baselines
Tech. Compilation@1 Pass@1 DSR@1 Repairable Ratio CodeBLEU Matchast
LLM-based with PE 57.9% 49.1% 69.6% 40.3% 0.604 0.401
Basic RAG 55.2% 48.3% 69.9% 41.8% 0.595 0.383
K3Trans 75.7% 67.7% 82.1% 44.6% 0.733 0.562
@@
Tech. Compilation@1 Pass@1 DSR@1 Repairable Ratio CodeBLEU Matchast
LLM-based with PE 57.9% 49.1% 69.6% 40.3% 0.604 0.401
@@
Table 3. Impact of Retrieval Strategy
Retrieval Strategy Compilation@1 Pass@1 DSR@1 Repairable Ratio CodeBLEU Matchast
BM25 + UniXcoder to get top1 64.8% 57.1% 74.7% 41.0% 0.697 0.531
BM25 + UniXcoder to get top2 68.0% 57.6% 72.8% 35.8% 0.706 0.536
BM25 + UniXcoder to get top3 67.2% 57.6% 73.6% 37.7% 0.704 0.533
As shown in Table 1, K3Trans achieved better results than two baselines in all execution-based metrics, with Compilation@1 and Pass@1 reaching 75.7% and 67.7%, respectively, substantially outperforming the best baseline LLM-based with PE by 17.8% and 18.6%. Significant improvements demonstrate the advantage of K3Trans in handling repository-level context code translation. Furthermore, K3Trans achieved 82.1% in DSR@1 and 44.6% in Repairable Rate, surpassing the baseline Basic RAG by 12.2% and 2.8%, respectively. This indicates that the code generated by K3Trans is of higher quality with higher repairability. In terms of match-based metrics, K3Trans also significantly outperforms the best baseline, achieving 0.733 of CodeBLEU and 0.562 of Matchast, representing substantial improvements over the baseline scores of 0.604 and 0.401. This fully demonstrates that the code generated by K3Trans is more similar to the ground truth and better captures the syntactic differences between programming languages.
In particular, LLM combined with the instructions designed in K3Trans outperforms Basic RAG adapted from previous work on four out of six metrics. This indicates that even the prompt design in K3Trans alone can effectively enhance LLM performance in repository-level context function translation.
{mdframed}
[linecolor=gray!25,roundcorner=12pt,backgroundcolor=myyellow!30,linewidth=3pt,innerleftmargin=2pt, leftmargin=0cm,rightmargin=0cm,topline=false,bottomline=false,rightline = false] Finding 1: K3Trans outperforms the baselines in all metrics. Even prompt design in K3Trans alone can achieve slightly better results than the baseline adapted from previous work.
@@ 4.2. RQ2: Ablation Study
K3Trans introduces three types of knowledge to improve LLM’s ability of code translation: 1) the successful translation function pair, i.e. previously successful translation knowledge, enhancing the LLM’s ability to recognize syntax differences between languages, 2) target-language code samples, i.e. retrieving similar target language code to lead LLM in generating code that conforms to the syntax of the target language, 3) dependency usage examples, i.e. providing relevant statement to teach LLM to correctly identify and invocate dependencies.
We explored the impact of different types of knowledge in the knowledge base on K3Trans by progressively reducing the provided knowledge categories. As shown in Table 2, each type of knowledge in the knowledge base significantly contributes to K3Trans ’s effectiveness in handling repository-level context code translation. With the inclusion of more types of knowledge, nearly every metric exhibits a clear trend of linear improvement. In particular, the comparison between LLM-based with PE and K1 demonstrates that dependency usage examples lead to a substantial increase in execution-based metrics, with a maximum relative improvement of 28.1%. This is because dependency usage examples provide the LLM with the correct invocation paths and methods for dependencies, effectively reducing the likelihood of LLMs incorrectly invoking dependencies, resulting in minimized compilation errors in the generated code, leading to more executable code.
@@
{mdframed}
[linecolor=gray!25,roundcorner=12pt,backgroundcolor=myyellow!30,linewidth=3pt,innerleftmargin=2pt, leftmargin=0cm,rightmargin=0cm,topline=false,bottomline=false,rightline = false] Finding 2: Each knowledge significantly contributes to K3Trans ’s effectiveness in handling repository-level context code translation, with dependency usage examples making the most notable contribution.
@@ 4.3. RQ3: Impact of Retrieval Strategy
Table 4. Translation performance under different stages of the knowledge base: 0%, 50%, and 100%.
Tech. Compilation@1 Pass@1 DSR@1 Repairable Ratio CodeBLEU Matchast
@@
100% knowledge base (K3Trans) 75.7% 67.7% 82.1% 44.6% 0.733 0.562

### v3

@@ 2.4. Knowledge-Augmented Code Translation
Figure 4. Prompt For Code Repair
LLM-based code repair. The code translated by LLMs may contain simple syntax errors (e.g., performing mutable operations on immutable variables). Previous work (Deligiannis et al., 2024) has shown that LLMs can repair certain errors in code using compiler error messages. After translation, K3Trans provides the LLM with specific instructions to leverage error message and iteratively refine translation results that fail testing due to compilation or functional errors. While this process can be conducted for multiple iterations, our implementation limits it to one iteration for cost-effectiveness. This approach further improves translation accuracy and assesses the percentage of erroneous translations that can be successfully repaired. The prompt used for code repair is shown as Fig. 4
@@ 3. Experiment Setup
  • RQ5 (Bad Case Study): Why does K3Trans fail in some translation tasks?
@@ 3.1. Benchmark
We choose RustRepoTrans (Ou et al., 2024) for the following reasons. RustRepoTrans is the first repository-level context code translation benchmark, constructed from GitHub projects that reflect the complexities of real-world software development. It consists of 375 function translation tasks at the repository level, where the source languages are C, Java, and Python, and the target language is Rust. Each task is accompanied by corresponding test cases, with an average test coverage exceeding 90%. As shown in previous work (Ou et al., 2024), LLMs exhibit much poorer performance on RustRepoTrans than on other benchmarks due to the repository-level dependencies and context. This shortcoming of LLM in dealing with real translation scenarios is exactly what K3Trans aims to solve.
Studied Programming Languages. We focus on repository-level context code translation for translating code from C, Java and Python to Rust, which are the same supported language pairs in RustRepoTrans. A more important reason is that Rust, as the target language, is much closer to the actual requirements of real-world code translation scenarios (Security, 2024). Furthermore, without language-specific adaptation, if K3Trans performs well with a low-resource language such as Rust, it can be reasonably inferred that K3Trans would also be applicable to other high-resource languages.
@@ 3.2. Baselines
We evaluate the performance of K3Trans with three baselines covering both LLM-based and rule-based methods: LLM-based methods include Basic RAG and LLM-based with prompt engineering (PE), while the rule-based method is represented by C2Rust.
@@
C2Rust (immunant, 2025), the predominant automated tool for converting C code to Rust, operates exclusively at the syntax level. It leverages the fact that Rust supports both unsafe constructs—which align closely with those in C—as well as features that enforce memory safety to translate each feature used by the C code into a possibly unsafe equivalent feature of Rust to resolve the syntactic discrepancies between C and Rust.
Table 1. Studied LLMs
Model Type Model Name Open-source reasoning Time Size General LLM Claude-3-5-Sonnet 2024.6 - DeepSeek-V3-0324 2025.3 671B DeepSeek-R1 2025.1 671B Code LLM Qwen-2.5-coder-32B 2024.9 32B Qwen-2.5-coder-14B 2024.9 14B
@@ 3.3. Metrics
  • Matchast: The Matchast is part of CodeBLEU, the reason we extracted it separately is to evaluate the ability of LLMs to recognize syntax differences. If the model can accurately identify syntactic differences and avoid translating check statements that are required in the source language but unnecessary in the target language (e.g. null pointer checks in C, which do not need to be explicitly performed in Rust), the generated code paths will be more similar to the ground truth, thus resulting in a higher Matchast.
Execution-based Metrics. The execution-based metrics aim to assess the quality of generated code through the execution status of the code. We employed Compilation@k, Pass@k, DSR@k and Repairable Rate for evaluation:
@@
    (4) $\text{Pass@k}=\mathbb{E}_{\text{Problems}}\left[1-\frac{\binom{n-c}{k}}{\binom{n}{k}}\right$]
@@
    (5) $\text{DSR@k}=\frac{1}{N}\sum_{i=1}^{N}S(i,k)$
  • Repairable Ratio: Repairable Ratio (RR), as expressed in Equation 6, is an evaluation metric for code translation tasks that measures the proportion of incorrect translations that can be automatically repaired using an LLM. The higher the value of this metric, the fewer interventions are needed to fix the code generated by the model, while initially incorrect, indicating higher quality of the generated code.
    (6) $\text{RR}=\frac{\text{DSR@1 - Pass@1}}{\text{1 - Pass@1}}$
@@ 3.4. Implementations
We construct the 100% state of the knowledge base, where all functions except the target function have been successfully translated and can serve as sources for dependency usage examples and translation function pair examples, along with all currently available GitHub open-source repositories to build target-language code samples. This setup, referred to as K3Trans in this paper, is used to evaluate the theoretical best performance that an LLM enhanced with self-evolving RAG can achieve in code translation tasks. As shown in Table 1, the models we select as the foundational models for K3Trans include general LLM and code LLM, open-source and closed-source models, reasoning models and non-reasoning models, as well as models with varying parameter sizes within the same series. This comprehensive selection demonstrates that K3Trans is applicable to various existing types of large models. However, due to resource constraints, after validating the generalizability of K3Trans in RQ1, we selected the best-performing model, Claude-3-5-Sonnet, as the representative model for further evaluation and analysis in the subsequent RQs. For the knowledge retrieval process, we utilize Elasticsearch (Elasticsearch, 2025) as our search engine, which based on the Lucene library using BM25 as the default score function.
@@ 4.1. RQ1: Performance of K3Trans
Table 2. Comparison of K3Trans and LLM-based Baselines
Model Tech. Compilation@1 Pass@1 DSR@1 RR CodeBLEU Matchast Claude-3-5-Sonnet Basic RAG 55.2% 48.3% 69.9% 41.8% 0.595 0.383 PE 57.9% 49.1% 69.6% 40.3% 0.604 0.401 K3Trans 75.7% 67.7% 82.1% 44.6% 0.733 0.562 DeepSeek-R1 Basic RAG 57.3% 46.4% 64.1% 33.0% 0.588 0.375 PE 61.1% 52.3% 65.6% 27.9% 0.605 0.400 K3Trans 71.2% 61.1% 74.4% 34.2% 0.725 0.553 DeepSeek-V3-0324 Basic RAG 42.1% 35.2% 48.5% 20.5% 0.584 0.366 PE 51.7% 43.7% 51.5% 13.9% 0.616 0.415 K3Trans 74.4% 63.2% 71.5% 22.6% 0.711 0.529 Qwen-2.5-coder-32B Basic RAG 31.2% 27.2% 37.3% 13.9% 0.576 0.352 PE 36.3% 31.5% 37.1 % 8.2% 0.585 0.358 K3Trans 53.6% 47.5% 56.8% 17.7% 0.696 0.503 Qwen-2.5-coder-14B Basic RAG 22.9% 19.2% 33.3% 17.5% 0.515 0.278 PE 25.6% 20.5% 29.6% 11.4% 0.465 0.233 K3Trans 53.3% 45.3% 54.7% 17.2% 0.684 0.493
Comparison with LLM-based baselines. As shown in Table 2, K3Trans achieved better results than both LLM-based baselines among all execution-based metrics for all studied LLMs. The best-performing model was Claude-3-5-Sonnet, with Compilation@1 and Pass@1 reaching 75.7% and 67.7%, respectively, substantially surpassing the best baseline PE by 17.8% and 18.6%. Furthermore, K3Trans with Claude-3-5-Sonnet achieved 82.1% in DSR@1 and 44.6% in RR, surpassing the baseline Basic RAG by 12.2% and 2.8%, respectively. This indicates that the code generated by K3Trans is of higher quality with higher repairability. In terms of match-based metrics, K3Trans also significantly outperforms the best baseline, achieving 0.733 of CodeBLEU and 0.562 of Matchast, representing substantial improvements over the baseline scores of 0.604 and 0.401. This fully demonstrates that the code generated by K3Trans is more similar to the ground truth and better captures the syntactic differences between programming languages.
The most significant improvements were observed on Qwen-2.5-coder-14B, with relative increases of 108.2% in Compilation@1 and 121.0% in Pass@1. Additionally, CodeBLEU and Match scores improved by 47.1% and 111.59%, respectively. It should also be noted that the highest relative improvements of all metrics were all observed on Qwen-2.5-coder-14B, the worst-performing model. This indicates that K3Trans can effectively compensate for the base model’s deficiencies in code translation, enhancing its capability in complex code scenarios.
In particular, LLMs combined with the instructions designed in K3Trans outperform Basic RAG, which is adapted from previous work, in multiple metrics such as Pass@1 and CodeBLEU on all models. This indicates that even the prompt design in K3Trans alone can effectively enhance LLM performance in repository-level context function translation.
Comparison with Rule-based baseline. Since C2Rust only supports translating code from C to Rust, we extracted 145 tasks from RustRepoTrans where the source language is C and the target language is Rust, and compared the performance of K3Trans and C2Rust on these tasks. Additionally, since rule-based methods lack self-debugging capabilities, we disregarded the DSR@1 and RR metrics. As shown in Table 3, K3Trans also achieved better results than the rule-based baseline across all execution-based metrics. C2Rust’s Compilation@1 and Pass@1 were only 0.7% and 0.7% (i.e., successfully translating only one task). This is because rule-based methods operate at the syntax level and rely on one-to-one code mapping, making them unable to handle scenarios involving dependency differences. Among the failures, misinterpretations due to Function Dependencies Difference and Data Type Dependencies Difference accounted for as high as 95.2% and 77.9%, respectively. We will provide a more detailed analysis in Section 4.5. In terms of match-based metrics, the rule-based baseline also performed poorly, with CodeBLEU and Matchast only 0.277 and 0.034, respectively. This is because C2Rust performs syntax-level, one-to-one translation of the source language code, resulting in low similarity with the target language version of the code.
Finding 1: K3Trans outperforms all baselines across all metrics among all models. Even the prompt design in K3Trans alone achieves slightly better results than the LLM-based baseline adapted from previous work. Furthermore, the rule-based method fails to effectively handle incremental translation scenarios with repository-level context.
Table 3. Comparison of K3Trans and a Rule-Based Baseline on C-to-Rust Translation Tasks
Tech. Model Compilation@1 Pass@1 CodeBLEU Matchast C2Rust - 0.7% 0.7% 0.277 0.034 K3Trans Claude-3-5-Sonnet 73.8% 61.4% 0.683 0.490 DeepSeek-R1 76.6% 62.1% 0.686 0.486 DeepSeek-V3-0324 75.9% 57.9% 0.672 0.466 Qwen-2.5-coder-32B 46.2% 37.2% 0.650 0.423 Qwen-2.5-coder-14B 51.0% 39.3% 0.640 0.424
@@
Tech. Compilation@1 Pass@1 DSR@1 RR CodeBLEU Matchast
PE 57.9% 49.1% 69.6% 40.3% 0.604 0.401
@@ 4.2. RQ2: Ablation Study
K3Trans introduces three types of knowledge to improve LLM’s ability of code translation: 1) dependency usage examples, i.e. providing relevant statement to teach LLM to correctly identify and invocate dependencies, 2) target-language code samples, i.e. retrieving similar target language code to lead LLM in generating code that conforms to the syntax of the target language, 3) the successful translation function pair, i.e. previously successful translation knowledge, enhancing the LLM’s ability to recognize syntax differences between languages.
We explored the impact of different types of knowledge in the knowledge base on K3Trans with Claude-3-5-Sonnet by progressively reducing the provided knowledge categories. As shown in Table 4, each type of knowledge in the knowledge base significantly contributes to K3Trans ’s effectiveness in handling repository-level context code translation. With the inclusion of more types of knowledge, nearly every metric exhibits a clear trend of linear improvement. In particular, the comparison between LLM-based with PE and K1 demonstrates that dependency usage examples lead to a substantial increase in execution-based metrics, with a maximum relative improvement of 28.1%. This is because dependency usage examples provide the LLM with the correct invocation paths and methods for dependencies, effectively reducing the likelihood of LLMs incorrectly invoking dependencies, resulting in minimized compilation errors in the generated code, leading to more executable code. More fine-grained interpretation of the role of dependency usage examples with concrete instances are provided in Section 4.5.
@@
Finding 2: Each knowledge significantly contributes to K3Trans ’s effectiveness in handling repository-level context code translation, with dependency usage examples making the most notable contribution.
@@ 4.3. RQ3: Impact of Retrieval Strategy
Table 5. Impact of Retrieval Strategy. The retrieval method is BM25 combined with UniXcoder.
Retrieval Strategy Compilation@1 Pass@1 DSR@1 RR CodeBLEU Matchast
Top1 64.8% 57.1% 74.7% 41.0% 0.697 0.531
Top2 68.0% 57.6% 72.8% 35.8% 0.706 0.536
Top3 67.2% 57.6% 73.6% 37.7% 0.704 0.533
Table 6. Translation performance under different stages of the knowledge base: 0%, 50%, and 100%. 100% corresponds to K3Trans.
Tech. Compilation@1 Pass@1 DSR@1 RR CodeBLEU Matchast
@@
100% knowledge base 75.7% 67.7% 82.1% 44.6% 0.733 0.562
As shown in Table 5, increasing the number of target-language code samples from one to two leads to improvements across Compilation@1, Pass@1, CodeBLEU, and Matchast, with Compilation@1 increasing notably from 64.8% to 68.0%. This indicates that adding more target-language code samples initially has a positive effect on the LLM. However, when the number is increased to three, the metrics no longer show significant improvements, and some, such as Compilation@1, even decrease, suggesting that excessive target-language code samples can introduce noise. Furthermore, as shown in RQ2 (Section 4.2), target-language code samples primarily impact match-based metrics like CodeBLEU and Match Score. Therefore, K3Trans ultimately sets the number of target-language code samples to two, as achieving the best performance on Compilation@1, Pass@1, CodeBLEU, and Matchast.
Finding 3: The optimal number of target-language code samples to provide in the repository-context code translation scenario is two.
4.4. RQ4: Impact of Self-evolution
To evaluate the impact of the self-evolution process on K3Trans, we design three knowledge base stages—0%, 50%, and 100%—to simulate different stages of self-evolution. Here, 0% corresponds to the baseline (LLM-based with Prompt Engineering), and 100% corresponds to K3Trans. To comprehensively assess the practicality of the self-evolution process in different translation scenarios and to account for the randomness in the translation order of the target functions, we construct 50% knowledge base as follows: Dependency usage examples: For each target function, we randomly retain 50% of the dependency usage examples available at the 100% stage. The remaining 50% dependencies have their dependency usage examples set to empty, representing no reference is available in current project’s context. Target-language Code Samples: We randomly select 50% projects from the collected projects and extract functions to construct the samples. Successful translation function pairs: For each target function, we randomly select 50% of the successful translation function pairs available at the 100% stage. These selected pairs are then used as candidates for the top-1 successful translation function pair retrieval.
Finding 4: As the self-evolution process progresses, the knowledge base continuously enhances the LLM’s performance across various aspects of the repository-level code translation.
4.5. RQ5: Case Study
To fully understand the advantages and limitations of K3Trans, we further manually analyze the successful examples of K3Trans that the baselines fail, and the failure examples of K3Trans.
Table 7. The Number of Top3 Error Causes of Dependencies Misuse of K3Trans and LLM-based Baseline
Model Tech. Error type Function Differences Misinterpretation ↓ Data Type Differences Misinterpretation ↓ Variable Differences Misinterpretation ↓ Claude-3-5-Sonnet Basic RAG 368 2 131 PE 191 0 58 K3Trans 247 0 40 DeepSeek-R1 Basic RAG 228 1 48 PE 172 0 32 K3Trans 256 2 24 DeepSeek-V3-0324 Basic RAG 420 1 188 PE 368 0 100 K3Trans 338 0 62 Qwen-2.5-coder-32B Basic RAG 679 11 220 PE 677 7 162 K3Trans 447 2 83 Qwen-2.5-coder-14B Basic RAG 1,288 11 219 PE 1,400 9 137 K3Trans 524 0 64 Total Basic RAG 2,983 25 806 PE 2,809 17 489 K3Trans 1,812 4 273
4.5.1. Analysis of Successful Examples
Since LLM translation failures on RustRepoTrans are attributed to Dependencies Misuse errors in as high as 77.9% of cases (Ou et al., 2024), we statistically analyzed the top three categories of Dependencies Misuse errors in the translation results of K3Trans and all baselines: Function Differences Misinterpretation, Data Type Differences Misinterpretation, and Variable Differences Misinterpretation, recording the specific count for each category.
In terms of LLM-based baselines, as shown in Table 7, compared to all LLM-based baselines, K3Trans effectively reduced the number of errors almost across all categories for all studied LLMs. Specifically, compared to the best-performing baseline PE, K3Trans demonstrated a reduction of 35.5% to 76.5% across all error categories. This indicates that K3Trans can significantly enhance the model’s ability to accurately identify and handle non one-to-one mapping dependency differences in incremental translation scenarios.
Additionally, compared to PE, Basic RAG showed similar or even higher error counts across these three dependency error categories for all studied LLMs. This suggests that even the prompt design in K3Trans (i.e., the PE baseline) alone slightly improves the model’s capability to handle repository-level context compared to the Basic RAG baseline, which is consistent with the conclusions in Section 4.1.
In terms of the Rule-based baseline, as shown in Table 8, C2Rust exhibited 3,307 Function Difference Misinterpretation errors across 138 out of 145 tasks, and 637 Data Type Differences Misinterpretation errors across 113 tasks, affecting 95.2% and 77.9% of the tasks, respectively. The reason is that rule-based methods operate at the syntax level and rely on one-to-one code mapping, making them unable to handle scenarios involving dependency differences. Specifically, as shown in Fig. 5, the Rule-based baseline performs a direct one-to-one translation of dependency call statements without any adaptation, failing to handle differences in repository-level context. However, C2Rust demonstrated the best performance in Variable Differences Misinterpretation. This is because C2Rust translates non-local variables by statically parsing the specific values of variables in the source language and directly using these values instead of variable names in the target language. While this approach effectively reduces compilation errors potentially caused by variable differences, it significantly compromises code readability and does not actually address repository-level differences between the source and target language versions.
Table 8. The Number of Top3 Error Causes of Dependencies Misuse of K3Trans and Rule-based Baseline
Tech. Model Error type Function Differences Misinterpretation ↓ Data Type Differences Misinterpretation ↓ Variable Differences Misinterpretation ↓ C2Rust - 3,307 637 26 K3Trans Claude-3-5-Sonnet 26 0 48 DeepSeek-R1 23 0 31 DeepSeek-V3-0324 35 0 34 Qwen-2.5-coder-32B 112 0 145 Qwen-2.5-coder-14B 129 0 127
Figure 5. Failure Example From Rule-based baseline
4.5.2. Analysis of Failure Examples
In particular, we separately analyze the proportions of compilation errors and functional implementation errors in bad cases both before and after self-debugging. Furthermore, we classify compilation errors into finer-grained categories based on the taxonomy defined in RustRepoTrans (Ou et al., 2024). This analysis helps us identify which errors persist despite the triple-knowledge enhancement, as well as which errors can be easily fixed through self-debugging and which cannot. By doing so, we can pinpoint areas where future work should focus on improving LLM capabilities to further enhance performance in repository-level context code translation.